# The Bayesian Separation Principle for Data-driven Control

### v1

@@ IV Separation principle for identification and control
Again, it is fundamental to emphasize that the existence or non-existence of this principle will depend on the context in which we work, the stated assumptions and also the significance of optimality. In particular
for the control problem, optimality will henceforth be understood as the definition of the best possible estimator of the MPC cost according to some useful metric, as the mean squared error. Hence, we say that the control input $u$ is optimal if it minimizes the optimal estimate of the MPC cost.
IV-A Fisherian setting
The mean squared error introduced in (3) measures the performance of the estimator $\hat{\theta}$, function of the identification data $\mathcal{D}$, in reconstructing $\theta$. Such notion immediately extends to measuring the distance between any function $g(\theta)$ and $g(\hat{\theta})$ as follows
@@ IV-A Fisherian setting
Also recalling its dependence on the unknown $\theta$, it is well known that it is not possible to obtain an optimal estimator in terms of MSE considering the class of all possible functions of the data. The simplest example is the constant estimator which always returns the same estimate $\bar{\theta}$ independently of the measurements in $\mathcal{D}$. It will outperform any other estimator if the true parameter vector is equal to $\bar{\theta}$ but its MSE can rapidly increase moving to other regions of the parameter space. For MPC, where the objective to estimate is $J(u;\theta)$, this means that it is not even possible to define the optimal solution adopting the direct approach to the control. A fortiori, we conclude that in the Fisherian setting a separation principle cannot exist using the MSE as optimality criterion if we work inside the class of all the possible estimators.
A separation principle instead arises if we restrict the solutions to the class of unbiased estimators and reason on an asymptotic basis. This principle can be derived from the theory of maximum likelihood (ML) estimators. In particular, we have already stressed that PEM concurs with the maximum likelihood estimator of $\theta$, here denoted by $\hat{\theta}^{ML}$, assuming Gaussian noise. Given any $g(\theta)$, the invariance principle then states that its ML estimator is obtained just replacing $\theta$ with $\hat{\theta}^{ML}$. This already hints at why the MPC solution obtained by exploiting this (preliminary) system identification step is asymptotically optimal. In fact, we can carry out such replacement in the ideal Fisherian objective (III) obtaining $J(u;\hat{\theta}^{ML})$ which so represents the maximum likelihood estimator of the control cost. Hence, if the postulated structure contains the true system $\mathcal{S}$, as the data size goes to infinity $J(u;\hat{\theta}^{ML})$ becomes optimal inside the class of unbiased estimators. Its mean converges to $J(u;\theta)$ for any $\theta$ and no other unbiased estimator of (III) can asymptotically have smaller variance. This argument still applies if more than one structure is postulated and BIC is used to control model complexity in (II) since we have seen that the exact dimension of $\theta$ will be returned asymptotically.
Of course, the limitation of this result is its asymptotical nature. An important point for real applications (that will come naturally in the Bayesian context) is to take into account also the uncertainty around the nominal model, and the MPC cost, defined by $\hat{\theta}^{ML}$. This aspect of the problem also gave rise to the rich literature on robust identification. It aims to derive confidence bounds useful e.g. to design robust controllers, capable of stabilizing a whole set of systems that contains $\mathcal{S}$ with high confidence, see e.g. [23, 30, 4].
@@
The separation principle here illustrated should not be misunderstood with the concept of sufficient statistic also discussed in [12]. Given a parametric model function of $\theta$, a sufficient statistic $T(\mathcal{D})$ is a transformation of the data set able to retain all the information contained in $\mathcal{D}$ for estimating $\theta$. This manifests itself in the fact that the MSE performance of any estimator of $g(\theta)$, which is modified by conditioning it on $T(\mathcal{D})$, improves or remains the same for any $\theta$. It follows that, if the minimum variance unbiased estimator of the MPC cost exists, it can be written as function only of $T(\mathcal{D})$. So, the construction of a sufficient statistic provides optimal compression of data (collected e.g. during an identification experiment) that can be used to estimate (III). However, the mere construction of $T(\mathcal{D})$ does not define or even guarantee the existence of the optimal estimator. Hence, it does not lead to fulfilling a separation principle according to the definition given above.
IV-B Bayesian parametric setting
Importantly, we will see that the separation principle applies in the Bayesian setting regardless of the data set size. This is first discussed in a parametric scenario where $\theta$ is still finite-dimensional but now represents a random vector.
Our assumptions (that will be relaxed later on) is that the system that generates the output data is still defined by a parametric structure $f$. But now it depends on a random (in place of deterministic) vector $\theta$ of known probability density function (pdf) $p(\theta)$. The stochastic system which generates the data is
$y(t)=f(y(t-1),y(t-2),\ldots,u^{t};\theta)+e(t)\$ (6)
where $u^{t}$ are the inputs up to instant $t$ while $e(t)$ is white noise, independent of $\theta$, of known pdf $p(e)$. In this Bayesian setting, the solution of the identification problem has to be seen as the entire posterior of $\theta$, calculated by the Bayes rule as follows
@@ IV-B Bayesian parametric setting
We use $\theta^{P}$ to indicate $\theta$ conditional on $\mathcal{D}$, whose pdf is indeed $p(\theta|\mathcal{D})$.
Consider now MPC in this stochastic setting. The first important issue is to obtain the counterpart of the Fisherian objective (III). The ideal control cost function to be minimized w.r.t. $u$ is now a stochastic variable, given by
$J_{u}(\theta,E)$ $=$ $\sum_{\ell=0}^{T-1}\|y_{\text{ref}}(t+\ell)$
-- $f(y(t+\ell-1),y(t+\ell-2),\ldots,u^{t+\ell};\theta)\|_{Q_{\ell}}^{2}$
$+$ $\sum_{\ell=0}^{T-1}\|u_{\text{ref}}(t+\ell)-u(t+\ell))\|_{R_{\ell}}^{2},$
whose randomness comes from the dependence on $\theta$ and on the future noises $e(t),e(t+1),\ldots$ contained in $E$. In fact, $y^{t-1}$ are not seen as stochastic, being set to the realizations of the corresponding past output random variables. Furthermore, the future outputs $y(t),y(t+1),\ldots$, which define recursively all the optimal predictors through
$\hat{y}(t+\ell)=f(y(t+\ell-1),y(t+\ell-2),\ldots,u^{t+\ell};\theta),$
are generated by $\theta$ and $e(t)$ through (6). Note that, differently from the Fisherian case, we can not already calculate the means of future outputs to obtain the predictors. In fact, this would remove the dependence on $\theta$ leading us to an objective far from the ideal one.
The bond between Bayesian identification and data-driven control comes about in a very natural way. Since $\theta$ and the future noises are not accessible to measurement, to obtain an estimator of (IV-B) first we condition the ideal cost $J_{u}$ on the identification data $\mathcal{D}$, obtaining the random variable $J_{u}|\mathcal{D}$. Then we take the expectation obtaining the estimator $\hat{J}_{u}$ function only of $\mathcal{D}$ (and the control input).
In view of the nature of $J_{u}$, since the future noises in $E$ are independent of $\mathcal{D}$, the conditional cost $J_{u}^{P}:=J_{u}|\mathcal{D}$ is function only of $\theta^{P}$ and $E$. Now it is useful to introduce the conditional outputs and predictors. Let
$\hat{y}^{P}(t):=f(y(t-1),y(t-2),\ldots,u^{t};\theta^{P})$
which also defines $y(t)$ conditional on $\mathcal{D}$ as follows
$y^{P}(t)=\hat{y}^{P}(t)+e(t).$
@@
$\hat{y}^{P}(t+1):=f(y^{P}(t),y(t-1),\ldots,u^{t+1};\theta^{P})$
so that
$y^{P}(t+1)=\hat{y}^{P}(t+1)+e(t+1)$
and so on. It comes that
$J^{P}_{u}$ $=$ $\sum_{\ell=0}^{T-1}\|y_{\text{ref}}(t+\ell)-\hat{y}^{P}(t+\ell)\|_{Q_{\ell}}^{2}$
$+$ $\sum_{\ell=0}^{T-1}\|u_{\text{ref}}(t+\ell)-u(t+\ell))\|_{R_{\ell}}^{2}$
@@
[…]
Bayesian estimation theory [1] ensures that this function $\hat{J_{u}}$ of the identification data $\mathcal{D}$ minimizes the mean squared error from $J_{u}(\theta,E)$ over the class of all possible estimators. Furthermore $\hat{J_{u}}$ is unbiased so that it is also the minimum variance estimator. This means that in the Bayesian scenario the question if indirect is better or not of direct has found a simple answer. Nothing can be better than computing the posterior and using it to average the uncertainty around the MPC objective. So,
@@
The cost (IV-B) could be also reformulated replacing the predictors $f(y(t+\ell-1),y(t+\ell-2),\ldots,u^{t+\ell};\theta)$ with the outputs $y(t+\ell)$. But nothing of substance would change. In fact, it is not difficult to see that the posterior mean $\hat{J_{u}}$ would only contain some additional irreducible error terms given by variances of the future noises in $E$.
To conclude, we observe that the cost (9) can be rewritten more explicitly in terms of the vector $Y^{P}_{\ell}:=[y^{P}(t),\dots,y^{P}(t+\ell-1)]^{\top}$, which is a function of $u(t),\dots,u(t+\ell-1)$, as follows
[…]
$+\sum_{\ell=0}^{T-1}\|u^{\text{ref}}_{t+\ell}-u_{t+\ell}\|^{2}_{R_{\ell}}$
[…]
$+\sum_{\ell=0}^{T-1}\|u^{\text{ref}}_{t+\ell}-u_{t+\ell}\|^{2}_{R_{\ell}}$
$=\sum_{\ell=0}^{T-1}\int\|y^{\text{ref}}_{t+\ell}-\mathbb{E}[\hat{y}^{P}_{t+\ell}|Y^{P}_{\ell}]\|^{2}_{Q_{\ell}}p(Y^{P}_{\ell})dY^{P}_{\ell}$
[…]
$+\!\sum_{\ell=0}^{T-1}\!\|u^{\text{ref}}_{t+\ell}\!-\!\hat{u}_{t+\ell}\|^{2}_{R_{\ell}}$ (10)
where we used the shorthand notation $y^{\text{ref}}_{t+\ell}:=y_{\text{ref}}(t+\ell)$, $u^{\text{ref}}_{t+\ell}:=u_{\text{ref}}(t+\ell)$, $\hat{y}^{P}_{t+\ell}:=\hat{y}^{P}(t+\ell)$, $u_{t+\ell}:=u(t+\ell)$. This reformulation will be exploited in the next section to derive more explicit and computationally tractable expressions of the MPC cost.
IV-C Bayesian nonparametric setting: Gaussian regression
The solution to the MPC problem derived in the previous setting could not yet be satisfactory under some aspects. We have seen that the dimension of $\theta$ is often unknown in real applications and its estimation is not trivial. Criteria such as BIC, very much in tune with this Bayesian setting, have also shown some weaknesses [33]. The formulation of a good finite-dimensional structure can therefore be difficult as well as the definition of a suitable prior on $\theta$. This kind of criticism is very much in line with the desire of direct approaches to break free from the model in some way. In the MPC literature many of the proposed techniques are called (or aspire to be) model free. This is somewhat utopian since in practice a data model, at least implicitly, needs always to be introduced. In statistical terms, the term model free should be connected with the use of nonparametric structures leading to null (or negligible) model bias. The latter has not to be confused with the estimation bias mentioned earlier. Specifically, we say that
the approach is model free or, equivalently, nonparametric if it has negligible model bias. This means that the system is sought in a space of such a large size (possibly even infinite-dimensional) that it practically contains all the possible systems and predictors of interest.
To obtain a Bayesian model free approach to MPC, our future objective is to cast Gaussian regression [40] in the control framework. First, our stochastic system $f$ in (6) is seen as the sum of a large (possibly infinite) number of (deterministic) basis functions $\phi_{k}$. Their stochastic coefficients are the components of $\theta$, modeled as zero-mean independent Gaussian random variables of variance $\lambda_{k}$, i.e.
$\theta_{k}\sim\mathcal{N}(0,\lambda_{k}).$ (11)
@@ IV-C Bayesian nonparametric setting: Gaussian regression
$x=[y(t-1)\ y(t-2)\ \ldots\ u^{t}]^{\top}$
so that we can write
@@
where $p$ could also grow to infinity. To simplify notation, assume that $f$ is real-valued and, just for a while, also that $x$ is scalar. This permits to introduce some simple and illustrative basis functions like sinusoids
@@
where $\eta$ is called kernel width. Combinations of these $\phi_{k}$ can approximate arbitrarily well any continuous function on any compact. However, it is not possible to learn a very large/infinite number of coefficients from the finite number of data contained in $\mathcal{D}$. But the prior (11) on the coefficients can address ill-posedness just introducing some high-level information about the system. In particular, kernel-based (Gaussian) regression is obtained by specifying how fast the variances $\lambda_{k}$ decay to zero so that
@@
Fulfillment of the second point leads to important computational advantages. No basis functions need to be stored in memory being all implicitly embedded in $K$, a point related to the so called kernel trick [42]. It allows to efficiently solve a variety of non-linear optimization problems that arise in regression and classification.
Using the sine waves (13) and the choice
@@
leads to the first-order spline kernel
@@
which includes just information on continuity of $f$ [47]. Increasing the value of $\alpha$ in (16) augments expected smoothness leading e.g. to the second-order spline kernel [5]. These two kernels are likely the most famous to estimate functions over one-dimensional domains.
Adopting (15) and further increasing the decay rate through
@@
leads to the widely used Gaussian kernel [44, 32]. It is infinitely differentiable and suited to describe functions particularly smooth. In the general multivariate case, where $x$ and $a$ can now be vectors containing past inputs and outputs, it assumes the form
@@
In the modeling of dynamic systems this permits to include the concept of fading memory. One can in fact partition $D$ into two diagonal blocks $D_{y}$ and $D_{u}$, one dedicated to outputs, the other to inputs, whose components tend to zero. In this way, the information that past inputs and outputs influence on the current output vanishes over time is included in the kernel.
@@
$Y=F\theta+E\$ (22)
where $Y$ is the vector collecting all the output data in $\mathcal{D}$, $F$ has two Toeplitz matrices side by side with past input and output data and $\theta$ now indicates the (column) random vector with the predictor impulse responses coefficients. In this case, a linear kernel can be used
@@
and the matrix $M$ represents the covariance of $\theta$. Similarly to what discussed about the matrix $D$ entering the nonlinear kernel (20), $M$ can be partitioned into two blocks $M_{y}$ and $M_{u}$, dedicated to outputs and inputs. Information on smooth exponential decay can then be introduced adopting the class of stable, TC or DC kernels reviewed in [33, 37]. For instance, the $(i,j)$ entry of $M_{y}$ (or of $M_{u}$) using TC is
[…]
@@
These example kernels thus provide nonparametric (model free) approaches where the system is described by a zero-mean Gaussian random field of covariance $K$:
@@
with $\lambda$ a positive scalar. It is now necessary to study the shape of the MPC cost resulting from such a modeling choice.
V MPC cost using Gaussian regression
Even if a Gaussian nonparametric (model free) technique can be adopted, another issue that could weaken the Bayesian separation principle is that direct approaches also wish to provide control objectives of simple structure, possibly depending on few parameters that can be tuned from data. Given a control input $u$, the evaluation of the Bayesian control objective $\hat{J_{u}}$ instead requires the solution of the integral (9) whose computational cost could be high, or even prohibitive for real time control applications. An important point is now to study $\hat{J_{u}}$ within the Gaussian regression framework, also understanding what parameters it depends on and how they can be estimated from the data. In what follows, just to simplify exposition, we still assume that the system outputs are scalar, so that $f$ is a real-valued Gaussian random field.
V-A Inference using Gaussian regression
According to the Bayesian separation principle, first we need to obtain the posterior of $f$ conditional on the identification data. These latter are contained in $\mathcal{D}$ and it is now convenient to denote them by
$z^{t}=[y(t-1)\ y(t-2)\ \ldots\ u^{t}]^{\top},\quad t=1,\ldots,N.$ (25)
Recall that data are generated according to (6) with $f$ that now satisfies (24) and is independent of the noises $e(t)$ which form a white Gaussian noise of variance $\sigma^{2}$. In the next proposition $z^{*}$ can be interpreted as a future input location where $f$ needs to be predicted to evaluate the MPC cost. The formulas given may seem standard, coming from basic results regarding jointly Gaussian vectors estimation [1]. Actually, there are some proof-related subtleties discussed in Appendix related to the fact that output data are also part of the input locations.

### v2

@@ 4 Separation principle for identification and control
Again, it is important to emphasize that the existence of this principle depends on the context in which we work, the stated assumptions, and also the definition of optimality. In particular
for the control problem, optimality will henceforth be understood as the definition of the best possible estimator of the MPC cost according to some useful metric, e.g. the mean squared error. Hence, we say that the control input $u$ is optimal if it minimizes the optimal estimate of the MPC cost.
4.1 Fisherian setting
[…]
@@ 4.1 Fisherian setting
Remembering its dependence on the unknown $\theta$, it is well known that it is not possible to obtain an optimal estimator in terms of MSE considering all possible functions of the data. The simplest example is the constant estimator which always returns the same estimate $\bar{\theta}$ independently of the measurements in $\mathcal{D}$. It will outperform any other estimator if the true parameter vector is equal to $\bar{\theta}$ but its MSE can rapidly increase moving to other regions of the parameter space. For MPC, where the objective to estimate is $J(u;\theta)$, this means that it is not even possible to define the optimal solution adopting the direct approach to the control. A fortiori, we conclude that in the Fisherian setting a separation principle cannot exist using the MSE as optimality criterion if we work inside the class of all the possible estimators.
A separation principle instead arises if we restrict the solutions to the class of unbiased estimators and reason on an asymptotic basis. It is derived from the theory of maximum likelihood (ML) estimators. In particular, we have already stressed that PEM concurs with the ML estimator of $\theta$, here denoted by $\hat{\theta}^{ML}$, assuming e.g. Gaussian noises and quadratic losses as done in (2). Given any $g(\theta)$, the invariance principle then states that its ML estimator is obtained just replacing $\theta$ with $\hat{\theta}^{ML}$. This already hints at why the MPC solution obtained by exploiting this (preliminary) system identification step is asymptotically optimal. In fact, we can carry out such replacement in the ideal Fisherian objective (5) obtaining $J(u;\hat{\theta}^{ML})$ which thus represents the maximum likelihood estimator of the control cost. Hence, if the postulated structure contains the true system $\mathcal{S}$, as the data size goes to infinity $J(u;\hat{\theta}^{ML})$ becomes optimal inside the class of unbiased estimators. Its mean converges to $J(u;\theta)$ for any $\theta$ and no other unbiased estimator of (5) can asymptotically have smaller variance. This argument still holds if more than one structure is postulated and BIC is used to control model complexity in (2) since we have seen that the exact dimension of $\theta$ is asymptotically returned. Of course, the limitation of this result relies in its asymptotic nature.
An important point for real applications (that will come naturally in the Bayesian context) is to take into account also the uncertainty around the nominal model $\hat{\theta}^{ML}$ and then around the MPC cost. This aspect of the problem also gave rise to the rich literature on robust identification. It aims to derive confidence bounds useful e.g. to design robust controllers, capable of stabilizing a whole set of systems that contains $\mathcal{S}$ with high confidence, see e.g. [17, 24, 2].
@@
The separation principle here illustrated should not be misunderstood with the concept of sufficient statistic also discussed in [9]. Given a model parametrized by $\theta$, a sufficient statistic $T(\mathcal{D})$ is a transformation of the data set $\mathcal{D}$ able to retain all the information contained in $\mathcal{D}$ for estimating $\theta$. This manifests itself in the fact that the MSE performance of any estimator of $g(\theta)$, which is modified by conditioning it on $T(\mathcal{D})$, improves or remains the same for any $\theta$. It follows that, if the minimum variance unbiased estimator of the MPC cost exists, it can be written as function only of $T(\mathcal{D})$. So, the construction of a sufficient statistic provides optimal compression of the data to estimate (5). However, the mere construction of $T(\mathcal{D})$ does not define or even guarantee the existence of the optimal estimator. Hence, it does not lead to fulfilling a separation principle according to the definition given above.
4.2 Bayesian parametric setting
We will now see that the separation principle applies in the Bayesian setting regardless of the data set size. This is first discussed in a parametric scenario where $\theta$ is still finite dimensional but now represents a random vector.
Our assumption (which will be relaxed later) is that the system can still be described by a parametric structure $f$. But now it depends on the random (instead of deterministic) vector $\theta$ of known probability density function (pdf) $p(\theta)$. Output data are generated as follows
$y(t)=f(y^{t-1},u^{t};\theta)+e(t)\\$ (6)
where $e(t)$ is white noise, independent of $\theta$, whose pdf is known. In this Bayesian setting, the solution of the identification problem has to be seen as the entire posterior of $\theta$, calculated by the Bayes rule as follows
@@ 4.2 Bayesian parametric setting
In the following we use $\theta^{P}$ to denote $\theta$ conditional on $\mathcal{D}$, whose pdf is therefore given by $p(\theta|\mathcal{D})$.
Let us consider MPC in this Bayesian setting. The first important issue is to obtain the counterpart of the Fisherian objective (5). The ideal control cost function to be minimized w.r.t. $u(\tau),u(\tau+1),\ldots$ is now a stochastic variable, given by
$J_{u}(\theta,E)$ $=\sum_{\ell=0}^{T-1}\|y_{\text{ref}}(\tau+\ell)-f(y^{\tau+\ell-1},u^{\tau+\ell};\theta)\|_{Q_{\ell}}^{2}$
$+\sum_{\ell=0}^{T-1}\|u_{\text{ref}}(\tau+\ell)-u(\tau+\ell))\|_{R_{\ell}}^{2},$ (7)
whose randomness comes from the dependence on $\theta$ and on the future noises $e(\tau),e(\tau+1),\ldots$ contained in $E$. In fact, the outputs in $y^{\tau-1}$ have not to be seen as stochastic: they have been already observed and set to their realizations. The future outputs $y(\tau),y(\tau+1),\ldots$ are instead stochastic and are generated by $\theta$ and the future noises through (6). Together with $y^{\tau-1}$, they define recursively all the optimal predictors through
$\hat{y}(\tau+\ell)=f(y^{\tau+\ell-1},u^{\tau+\ell};\theta).$
Note that, differently from the Fisherian case, the predictors depend explicitly on the future outputs (and are therefore random variables). They are not defined by computing their means because this would remove the dependence on $\theta$ leading us to a control objective far from the ideal one.
The bond between Bayesian identification and data-driven control comes about in a very natural way. Since $\theta$ and the future noises are not accessible to measurement, to obtain an estimator of (4.2) first we condition $J_{u}$ on the identification data $\mathcal{D}$, obtaining the following key random variable
$J_{u}^{P}:=J_{u}\,|\,\mathcal{D}$
which corresponds to the conditional cost. Then we take its expectation obtaining the estimator $\hat{J}_{u}$ function only of $\mathcal{D}$ (and the control input).
Since the future noises in $E$ are independent of $\mathcal{D}$, the conditional cost $J_{u}^{P}$ actually depends only on $\theta^{P}$ and $E$. Let us define the conditional predictor at instant $\tau$ as follows
$\hat{y}^{P}(\tau):=f(y^{\tau-1},u^{\tau};\theta^{P})$
so that $y(\tau)$ conditional on $\mathcal{D}$ is
$y^{P}(\tau):=\hat{y}^{P}(\tau)+e(\tau).$
@@
$\hat{y}^{P}(\tau+1):=f(y^{P}(\tau),y^{\tau-1},u^{\tau+1};\theta^{P})$
so that the next conditional output is
$y^{P}(\tau+1):=\hat{y}^{P}(\tau+1)+e(\tau+1)$
and so on. We can now write the conditional MPC cost as
$J^{P}_{u}$ $=\sum_{\ell=0}^{T-1}\|y_{\text{ref}}(\tau+\ell)-\hat{y}^{P}(\tau+\ell)\|_{Q_{\ell}}^{2}$
$+\sum_{\ell=0}^{T-1}\|u_{\text{ref}}(\tau+\ell)-u(\tau+\ell))\|_{R_{\ell}}^{2}$ (8)
@@
$\hat{J_{u}}=\int J^{P}_{u}(\theta,E)p(\theta|\mathcal{D})p(E)d\theta dE.$ (9)
Bayesian estimation theory [1] ensures that this function $\hat{J_{u}}$ of the identification data $\mathcal{D}$ minimizes the mean squared error from $J_{u}(\theta,E)$ over the class of all possible estimators. Furthermore, $\hat{J_{u}}$ is unbiased and gives the minimum variance estimate of the MPC cost. This means that in the Bayesian scenario the question if indirect is better or not than direct has found a simple answer: nothing can be better than computing the posterior, using it to average the uncertainty around the MPC objective and obtaining the control input as the minimizer of the resulting estimate $\hat{J_{u}}$. So,
@@
The cost (4.2) could be also reformulated replacing each predictor $f(y^{\tau+\ell-1},u^{\tau+\ell};\theta)$ with the output $y(\tau+\ell)$. But nothing substantial would change: the posterior mean $\hat{J_{u}}$ would just contain some additional irreducible error terms given by variances of the future noises in $E$.
4.3 Bayesian nonparametric setting: Gaussian regression
The solution to the MPC problem derived in the previous setting could not yet be satisfactory under some aspects. We have seen that the dimension of $\theta$ is often unknown in real applications and its estimation is not trivial. Criteria such as AIC or BIC, very much in tune with this Bayesian setting, have also shown some weaknesses to control complexity [30]. The formulation of a finite-dimensional structure with good prediction capability can therefore be difficult as well as the definition of a suitable prior on $\theta$. This kind of criticism is very much in line with the desire of direct approaches to break free from the model in some way. In the MPC literature many of the proposed techniques are called (or aspire to be) model-free. This is somewhat utopian since in practice a data model needs always to be introduced, at least implicitly. In statistical terms, the term model-free should be connected with the use of nonparametric structures leading to null (or negligible) model bias. The latter has not to be confused with the estimation bias mentioned earlier. Specifically, we say that
the approach is model-free or, equivalently, nonparametric if it has negligible model bias. This means that the system is searched in a space of such large size (possibly even infinite dimensional) that it contains virtually all possible systems and predictors of interest.
One way to get a Bayesian model-free approach to MPC is to cast Gaussian regression [37] into the control framework. There are many different ways of introducing this estimation technique. For example, Neal in [28] considered a neural network with independent and random weights and showed that the input-output map becomes a Gaussian process as the width of the network grows to infinity. This construction illustrates well the richness that this type of model can have. Similarly, below we introduce Gaussian regression starting from a wide range of basis functions whose coefficients are random variables.
First, our stochastic system $f$ in (6) is seen as the sum of a large, possibly infinite, number of (deterministic) basis functions $\phi_{k}$. Their stochastic coefficients are the components of $\theta$, modeled as zero-mean independent Gaussian random variables of variance $\zeta_{k}$, i.e.
$\theta_{k}\sim\mathcal{N}(0,\zeta_{k}).$ (10)
@@ 4.3 Bayesian nonparametric setting: Gaussian regression
$x=[y^{t-1}\ \ u^{t}]^{\top}$
where, in the case of NARX models, $y^{t-1}$ and $u^{t}$ going back in time only to the instant tmt-m. Now, we can write
@@
where $p$ could also grow to infinity. To further simplify notation, we also assume that $f$ is real-valued and, just for a while, that $x$ is scalar. This permits to introduce some simple and illustrative basis functions like sinusoids
@@
where $\eta$ is called kernel width. A common feature of these $\phi_{k}$ is that they are universal, i.e., they are able to approximate arbitrarily well any continuous function on any compact set. However, it is not possible to learn a very large/infinite number of coefficients from the finite number of data contained in $\mathcal{D}$. This intrinsic ill-posedness is overcome by the prior (10), whose aim is to specify some high-level information about the system. In particular, kernel-based (Gaussian) regression is obtained by specifying how quickly the variances $\zeta_{k}$ decay to zero, so that
  • 1.
@@
Fulfillment of the second point leads to important computational advantages. No basis functions need to be stored in memory being all implicitly embedded in $K$, a point related to the so-called kernel trick [39]. It allows the efficient solution of a wide variety of nonlinear optimization problems for regression and classification.
Using the sine waves (12), the choice of the variances in (10) given by
@@
defines the first-order spline kernel
@@
which includes just information on continuity of $f$ [43]. Increasing the value of $\alpha$ in (15) augments expected smoothness leading, e.g., to the second-order spline kernel [3]. These two covariances are likely the most famous to estimate functions over one-dimensional domains.
Adopting (14) and further increasing the decay rate of the variances as follows
@@
defines the widely used Gaussian kernel [41, 27]. It is infinitely differentiable and suited to describe functions particularly smooth. In the general multivariate case, where $x$ and $a$ can now be vectors containing past inputs and outputs, it assumes the form
@@
This leads to a form of physics-informed machine learning which makes use of regularizers which incorporate information on dynamical systems. In particular, the concept of fading memory can be included by partitioning $D$ into two diagonal blocks $D_{y}$ and $D_{u}$ one dedicated to outputs, the other to inputs, whose components tend to zero. For example, the entries along the diagonal of $D_{y}$ (or of $D_{u}$) can be defined by
$D_{y}]_{ii}\propto\alpha^{i}\ \ \mbox{with}\ \ 0<\alpha<1$
to model exponential decay with the rate $\alpha$. The kernel thus takes into account the fact that the influence of past inputs and outputs on the current output disappears over time.
@@
$Y=F\theta+E\\$ (21)
where $Y$ is the vector collecting all the output data in $\mathcal{D}$, $F$ has two Toeplitz matrices side by side with past input and output data and $\theta$ now indicates the (column) random vector with the predictor impulse responses coefficients. In this case, one can use a linear kernel
@@
and one can easily see that the matrix $M$ now represents the covariance of $\theta$. Similarly to what discussed about the matrix $D$ entering the nonlinear kernel (19), $M$ can be partitioned into two blocks $M_{y}$ and $M_{u}$, dedicated to outputs and inputs. Information on smooth exponential decay can then be introduced adopting the class of stable, TC or DC kernels reviewed in [30, 34]. For instance, the $(i,j)$ entry of $M_{y}$ (or of $M_{u}$) using TC is
$M_{y}]_{i,j}=\alpha^{\max(i,j)},\ 0\leq\alpha<1$
@@
In this section we have thus seen that in the Bayesian setting $f$ is seen as a random surface and
kernels are a way to define a nonparametric (model-free) approach where the system is a zero-mean Gaussian random field of covariance $K$:
@@
with $\lambda$ being a positive scalar.
It is now necessary to study the shape of the MPC cost resulting from such a modeling choice.
5 MPC cost using Gaussian regression
An issue that could weaken the Bayesian separation principle is that direct approaches also wish to provide control objectives of simple structure, possibly depending on few parameters that can be tuned from data. Given a control input $u$, the evaluation of the Bayesian control objective $\hat{J_{u}}$ instead requires the solution of the integral (9) whose computational cost could be high, or even prohibitive for real-time control applications. Thus, it is now important to study the estimator $\hat{J_{u}}$ of the MPC cost reported in (9) within the Gaussian regression framework, also understanding what parameters it depends on and how they can be estimated from the data.
5.1 Inference using Gaussian regression
In what follows, just to simplify exposition, the system output is assumed scalar, so that $f$ is a priori described as a real-valued Gaussian random field. According to the Bayesian separation principle, first we need to obtain the posterior of $f$ conditional on the input-output measurements taken before the time $\tau$ when the control phase starts.
It is convenient to build from the identification data contained in $\mathcal{D}$ the input locations
$z^{t}=[y^{t-1}\ \ u^{t}]^{\top},\quad t=1,\ldots,N,$ (24)
a notation that makes them “dynamic” in light of the dependence on time tt [^3] Strictly speaking, these input locations contain also some past data not contained in $\mathcal{D}$, e.g. input and outputs measured before $t=1$. If such data are not available one can handle initial conditions effects as e.g. described in [23, Section 3.2].. Recall that the $z^{t}$ are generated according to (6) with $f$ that satisfies (23) and is independent of the noises $e(t)$ which form a white Gaussian noise of variance $\sigma^{2}$. In the next proposition $z^{*}$ can be interpreted as a future input location where $f$ needs to be predicted to evaluate the MPC cost. The formulas given may seem standard, coming from basic results regarding jointly Gaussian vectors estimation [1]. Actually, there are some proof-related subtleties discussed in Appendix related to the fact that output data are also part of the input locations.